Find two numbers given that their sum is 48 and the smaller number is equal to one fifth of the larger number

Answer by clearblueskai(22)

let x be the smaller number and y the larger number
x + y = 48
x = 1/5y
substitute the second equation to the first equation
(1/5y) + y = 48
6/5 y = 48
Multiply both sides by 5/6 to cancel the numerical coefficient 6/5
y = (48)(5/6)
y = 40
x + 40 = 48
x = 48 - 40
x = 8
smaller number is 8
larger number is 40